OpenStudy (anonymous):

How so you solve for C in this equation? a=b(1/c-1/d)

OpenStudy (anonymous):

Do you have any ideas?

OpenStudy (anonymous):

Well I was thinking that I had to distribute b to (1/c-1/d)

OpenStudy (anonymous):

Yes. Correct. Then what do you think you do?

OpenStudy (hba):

Well follow those steps. 1)divide b on both sides. 2)Take LCM I'll continue when you are done with those steps.

OpenStudy (anonymous):

is it cd?

OpenStudy (anonymous):

wait. I'm confused now. ahah

OpenStudy (anonymous):

Okay, First you want to distribute b to (1/c - 1/d) then you want to add b to both sides and that would equal 1/c. Then you would do c = 1/(a/b + 1/d) and then c = b*d/(ad +b).

OpenStudy (anonymous):

Ohhh, ok, that you so much!

OpenStudy (anonymous):

You're welcome! :)